OpenStudy (anonymous):

Find solutions to: x(dy/dx)=(x^2)*(e^-x) + y satisfying y(1) = 0

myininaya (myininaya):

it looks like we can write this in the form: \[y'+p(x)y=q(x)\] \[xy'=x^2e^{-x}+y =>xy'-y=x^2e^{-x}=> y'-\frac{1}{x}y=xe^{-x}\]

myininaya (myininaya):

now this is pretty easy to solve first thing we need to do is multiply both sides by \[v(x)=e^{\int\limits_{}^{}-\frac{1}{x} dx}=e^{-lnx}=\frac{1}{x}\] \[\frac{1}{x}y'-\frac{1}{x} \cdot \frac{1}{x} y=e^{-x}\] we did so we could write \[(\frac{1}{x}y)'=e^{-x}\]

myininaya (myininaya):

now we need to integrate both sides

myininaya (myininaya):

\[\frac{1}{x}y=-e^{-x}+C\]

myininaya (myininaya):

\[\frac{1}{1} (0)=-e^{-1}+C\] after pluggin' x=1 and y=0 we can solve for C

myininaya (myininaya):

\[C=e^{-1}=\frac{1}{e}\]

myininaya (myininaya):

so we have \[\frac{1}{x}y=-e^{-x}+\frac{1}{e}\]

OpenStudy (anonymous):

That is a good solution using the integration constant \[v(x) = e ^{\int\limits_{}^{}-p(x)dx}\] Additionally you can use an equation that is also derived from the integration constant for the general form to give you the solution.\[y \prime+p(x)y=q(x)\] \[y = \ \frac{1}{v(x)} \ \int\limits_{}^{}q(x)v(x)dx+C\] By the way the final step for your solution will be: \[y = C _{1}x-xe ^{-x}\]

OpenStudy (anonymous):

Okay yeah and the value of constant \[C _{1}=\ \frac{1}{e} \]

myininaya (myininaya):

formulas are too easy to use lol

OpenStudy (anonymous):

Yes you're right..I particularly don't go that route but I like it because my prof. did a great job deriving it from general formula and some people like using it too

OpenStudy (anonymous):

I don't mind using formulas that I can derive completely it makes more sense that way.

myininaya (myininaya):

i like doing it long way because it makes more sense to me why that is the answer